# RG approach to the inviscid limit for shell models of turbulence

### v1

@@ Abstract
We consider an initial value problem for shell models that mimic turbulent velocity fluctuations over a geometric sequence of scales. Our goal is to study the convergence of solutions in the inviscid (more generally, vanishing regularization) limit and explain the universality of both the limiting solutions and the convergence process. We develop a renormalization group (RG) formalism representing this limit as dynamics in a space of flow maps. For the dyadic shell model, the RG dynamics has a fixed-point attractor, which determines universal limiting solutions. Deviations from the limiting solutions are also universal and given by a leading eigenmode (eigenvalue and eigenvector) of the linearized RG operator. Application to the Gledzer and Sabra shell models reveal the respective quasiperiodic and chaotic RG dynamics.
@@ 1 Introduction
Our main result is the development of the renormalization group (RG) formalism. It describes the limit of vanishing regularization as the dynamics of an RG operator acting in the space of flow maps. A similar formalism was presented earlier for models on a discrete space-time lattice [35, 34]. The extension to shell models in which time is continuous is carried out by introducing the concept of canonical (symmetry preserving) regularizations. For the dyadic model, the RG approach not only justifies the existence and universality of the limiting solutions as a fixed-point attractor of the RG dynamics, but also predicts new universal properties. This new universality is related to the leading eigenmode (eigenvalue and eigenvector) of the linearized RG operator, which determine deviations of regularized solutions from their limiting form. The usual viscous regularization is not canonical (it violates time scale invariance), but its RG analysis is mediated by auxiliary canonical regularizations. We also apply the RG approach to the Gledzer and Sabra shell models. In the former case, we show that the RG dynamics has a quisiperiodic attractor governing the vanishing regularization limit. For the Sabra model, the RG dynamics is chaotic with double exponential divergence of solutions.
In general, an RG formalism explores self-similarity of a system manifested at a large range of scales, but its precise form may vary a lot depending on the problem. In our case, this is the self-similarity of equations of motion corresponding to the ideal system. As we showed in [34], our RG approach has much in common with the Feigenbaum theory [17]. In particular, one can establish the analytical similarity for the model of digital turbulence. However, our RG operator contain in its definition a specific form of large-scale (ideal) interactions. Opposite to the Wilson RG formalism [40] and its extension to fluid dynamics [4], we do not coarse-grain system properties but rather keep all details at small scales intact. One can also recognize in our formalism some features of the inverse RG [15, 20, 23], because our RG operator acts by adding an extra largest scale and reconstructing the ideal dynamics at that scale.
The paper is organized as follows. Section 2 introduces shell models. Section 3 defines the RG operator and relates it to a class of canonical regularizations. Section 4 studies the limit of vanishing regularization in terms of the fixed-point RG attractor for the dyadic model. Section 5 extends the results to the (non-canonical) viscous regularization. Section 6 studies the quasiperiodic RG attractor in the Gledzer shell model. Section 7 describes the chaotic RG dynamics in the Sabra model. Section 8 summarizes the results and discusses further developments. Some technical derivations are gathered in the Appendix.
@@ 2 Ideal and viscous dyadic models
where the tildes denote a new solution for new initial and boundary conditions. Similarly, the time scaling symmetry takes the form
$\tilde{u}_{n}(t)=\alpha u_{n}(\alpha t),\quad\tilde{a}_{n}=\alpha a_{n},\quad\tilde{b}(t)=\alpha b(\alpha t),$ (2.6)
@@ 3 Canonical regularizations and RG operator
Here the arguments are the initial conditions $a=(a_{n})_{n\geq 1}$ and the boundary function $b\in C^{1}$. In analogy with the dynamical systems theory, we call $\Phi$ a flow map of a regularized IBVP. The flow map is assumed to have the following properties. Every solution (3.1) must satisfy the respective initial and boundary conditions (2.3) and (2.4). The causality property requires that the value of $\Phi_{t}(a,b)$ at a given time depends on the boundary function $b$ only from the past time interval $[0,t]$. Finally, $\Phi$ satisfies the usual semigroup relation (adapted to the time-dependent boundary condition) as
$\Phi_{t^{\prime}+t}(a,b)=\Phi_{t^{\prime}}\big{(}\Phi_{t}(a,b),b_{t}\big{)},\quad t,t^{\prime}\geq 0,$ (3.2)
where $b_{t}(t^{\prime})=b(t+t^{\prime})$ is the time-shifted boundary function.
@@
A flow map $\Phi$ is called renormalizable if there exists a unique flow map denoted as $\mathcal{R}[\Phi]$ with the following property: for any initial and boundary conditions $a$ and $b$, the solution $u(t)=\mathcal{R}[\Phi]_{t}(a,b)$ satisfies the relations
$\frac{du_{1}}{dt}=k_{1}f_{1},\quad u_{n}(t)=\frac{\tilde{u}_{n-1}(t)}{\lambda},\quad n\geq 2,$ (3.3)
where $\tilde{u}(t)=\Phi_{t}(\tilde{a},\tilde{b})$ with $\tilde{a}$ and $\tilde{b}$ defined as
$\tilde{a}_{n}=\lambda a_{n+1},\ n\geq 1;\quad\tilde{b}(t)=\lambda u_{1}(t).$ (3.4)
The transformation $\Phi\mapsto\mathcal{R}[\Phi]$ is called the RG operator.
@@
If $\Phi$ is a renormalizable $N$-level regularization, then $\mathcal{R}[\Phi]$ is an $(N+1)$-level regularization.
@@
One can see that the second relation in Eq. (3.3) and expressions (3.4) coincide with the scaling relation (2.5). Since $\tilde{u}(t)$ is generated by the $N$-level regularization $\Phi$, it satisfies the ideal Eq. (2.1) for the shells $n=1,\ldots,N$. Then one can check that $u_{n}(t)$ satisfy the ideal Eq. (2.1) for the shells $n=2,\ldots,N+1$. The first relation in Eq. (3.3) provides the remaining ideal equation for $n=1$. ∎
For now, Definition 1 has an important flaw. Namely, combining the two symmetries (2.5) and (2.6), one can replace Eqs. (3.3) and (3.4) by
$\frac{du_{1}}{dt}=k_{1}f_{1},\quad u_{n}(t)=\frac{\tilde{u}_{n-1}(t/\alpha)}{\lambda\alpha},\quad n\geq 2,$ (3.5)
where $\tilde{u}(t)=\Phi_{t}(\tilde{a},\tilde{b})$ with $\tilde{a}_{n}=\lambda\alpha a_{n+1}$ and $\tilde{b}(t)=\lambda\alpha u_{1}(\alpha t)$. It is straightforward to check that Proposition 1 remains valid for this new formulation. This means that our definition of the RG operator is ambiguous. This ambiguity is eliminated by considering a specific subclass of canonical regularizations, as we show next.
@@
$\Phi_{t}(\tilde{a},\tilde{b})=\alpha\Phi_{\alpha t}(a,b),\quad\tilde{a}_{n}=\alpha a_{n},\quad\tilde{b}(t)=\alpha b(\alpha t),$ (3.6)
for any $a$, $b$ and $\alpha>0$.
@@
The important property of the RG operator $\mathcal{R}$ is that it is uniquely defined by the ideal model. Indeed, the ideal coupling function $f_{1}$ is the only model-dependent element in Definition 1. The space of canonical regularizations (the domain of $\mathcal{R}$) also depends on properties (symmetries) of the ideal model only. On the contrary, all the information on how solutions are regularized is contained in the flow map $\Phi$ itself. These features highlight the distinct roles of the ideal system and regularization in the RG dynamics: the first determines the RG operator $\mathcal{R}$, and the second the initial flow map $\Phi$.
@@
We finish this section with a few additional remarks. The viscous regularization (2.7) is not canonical, because the dissipative term is linear (not quadratic). This leads to important consequences that we investigate later in Section 5. The cutoff of models (3.7) mimics Large Eddy Simulation (LES) closures in fluid dynamics [37]: the dissipative term can be written as $k_{n}|u_{n}|u_{n}=\nu_{n}(u)k_{n}^{2}u_{n}$ with the effective eddy-viscosity $\nu_{n}(u)=|u_{n}|/k_{n}$. Finally, note that Eqs. (3.3) and (3.4) express the variables $u_{n}(t)$ for $n\geq 2$ in terms of the full history of $u_{1}$ in the time interval $[0,t]$. Hence, the first relation in Eq. (3.3) represents a functional (not just an ordinary differential) equation for $u_{1}(t)$.
@@ 4 Fixed-point RG attractor
An immediate consequence of the limit (4.2) is that the limiting flow map provides solutions $u^{\infty}(t)=\Phi^{\infty}_{t}(a,b)$ of the ideal IBVP. Moreover, these limiting solutions are universal: they do not depend on the choice of regularization, as long as the initial flow map $\Phi^{(0)}$ belongs to the basin of the attraction. Taking the limit $N\to\infty$ in both sides of Eq. (4.1), one can see that $\Phi^{\infty}$ satisfies the fixed-point condition $\Phi^{\infty}=\mathcal{R}[\Phi^{\infty}]$. This fixed-point condition is understood in the sense that Eqs. (3.3) and (3.4) of Definition 1 are satisfied for $\Phi=\Phi^{\infty}$ and $\mathcal{R}[\Phi]=\Phi^{\infty}$.
@@
We now verify both relations (4.2) and (4.7) numerically for the regularized systems (3.7). In numerical simulations, we consider three different canonical regularization with the parameters $J=1,2,3$. The respective RG iterations are given by Eqs. (3.8) and (4.7) as $\Phi^{(N,J)}\approx\Phi^{\infty}+c_{J}\rho^{N}\Omega$. In this expression, the dependence on the regularization model $(N,J)$ reduces to the single constant factor $c_{J}\rho^{N}$. For numerical simulations, we consider two different initial conditions
@@
Numerical simulations are performed using the ode45 and ode15s solvers in Matlab with very high accuracy.
@@ 5 RG approach to the viscous model
Figure 2: (a) Regularized IBVP solutions for different models with the initial condition $\mathrm{IC}_{2}$. Their collapse confirms the convergence to the universal solution of the ideal IBVP. (b) Normalized deviations $\delta u_{1}(t)$ for different regularization models. Collapse of the graphs for large $N$ confirms the universal eigenmode correction for the canonical models $(N,J)$ and $(N,\beta)$. For the viscous model, the normalized deviation (green line) has a different (non-universal) shape. (c,d) Dependence of $N_{t}^{(\nu)}(a,b)$ and $\beta_{t}^{(\nu)}(a,b)$ on time for decreasing viscosities $\nu$.
@@ 5.1 Auxiliary regularized model
Hence, the first-shell Eq. (5.3) takes the ideal form asymptotically for large $N$, i.e., it yields the first relation in Eq. (3.3) of Definition 1. The rest of the argument is the same as for the regularized system (3.7) (see Section 3 and Appendix 9.2) and leads to the relation
@@ 5.2 Viscous model and inviscid limit
$u^{(\nu)}(t+dt)=\Phi^{(\nu)}_{dt}\left(u^{(\nu)}(t),b_{t}\right)=\Phi^{(N,\beta)}_{dt}\left(u^{(\nu)}(t),b_{t}\right),\quad\beta=\frac{\nu k_{N}}{|u_{N}^{(\nu)}(t)|}.$ (5.9)
Let us select $N=N_{t}^{(\nu)}(a,b)$ as the largest shell number providing the parameter $\beta=\beta_{t}^{(\nu)}(a,b)=\nu k_{N}/|u_{N}^{(\nu)}(t)|\leq 1$. The time dependence of such $N$ and $\beta$ for different (decreasing) viscosities $\nu$ and specific initial and boundary conditions are shown Fig. 2(c,d). One can see that the inviscid limit $\nu\to 0$ implies $N\to\infty$. We argued in Section 5.1 that $\Phi^{(N,\beta)}\to\Phi^{\infty}$ as $N\to\infty$. Assuming that this convergence is uniform with respect to $\beta\leq 1$, the identity (5.9) implies that the evolution of the viscous solution can be approximated by the limiting flow map $\Phi^{\infty}$ at all times in the inviscid limit. This explains why the viscous regularization yields the same inviscid limit $\Phi^{(\nu)}\to\Phi^{\infty}$. This convergence is verified numerically in Fig. 2(a) by comparing solutions for $\nu=10^{-8}$ and $(N,\beta)=(20,1)$.
@@ 5.3 Asymptotic form of deviations
$u^{(\nu_{N})}(t+dt)\approx\Phi^{\infty}_{dt}\big{(}u^{(\nu_{N})}(t),b_{t}\big{)}+c_{\beta}\rho^{N}\Omega_{dt}\big{(}u^{(\nu_{N})}(t),b_{t}\big{)}.$ (5.11)
@@
$\Phi^{\infty}_{dt}\big{(}u^{(\nu_{N})}(t),b_{t}\big{)}\approx\Phi^{\infty}_{dt}\big{(}u^{\infty}(t),b_{t}\big{)}+\rho^{N}\delta\Phi^{\infty}_{dt}\big{(}v(t);u^{\infty}(t),b_{t}\big{)},$ (5.13)
where $\delta\Phi^{\infty}_{dt}(\delta u;u^{\infty}(t),b_{t})$ is the variational derivative of $\Phi^{\infty}_{dt}(u^{\infty}(t)+\delta u,b_{t})$. Similarly, the last term in Eq. (5.11) to the first-order becomes
$c_{\beta}\rho^{N}\Omega_{dt}\big{(}u^{(\nu_{N})}(t),b_{t}\big{)}\approx c_{\beta}\rho^{N}\Omega_{dt}\big{(}u^{\infty}(t),b_{t}\big{)}.$ (5.14)
Finally, using Eqs. (5.12)–(5.14) in the relation (5.11), cancelling the zero-oder terms $u^{\infty}(t+dt)=\Phi^{\infty}_{dt}\big{(}u^{\infty}(t),b_{t}\big{)}$ and then the common factor $\rho^{N}$, we obtain
$v(t+dt)\approx\delta\Phi^{\infty}_{dt}\left(v(t);u^{\infty}(t),b_{t}\right)+c_{\beta}\Omega_{dt}\big{(}u^{\infty}(t),b_{t}\big{)}.$ (5.15)
@@
6 Quasiperiodic RG attractor
In this section, we further explore the dynamical systems approach to the vanishing regularization limit. We present here a more complex example of a quasiperiodic RG attractor. In this case regularized solutions do not converge in the limit $N\to\infty$, but one can describe them universally in terms of a one-parameter family of flow maps.

### v2

@@ Abstract
We consider an initial value problem for shell models that mimic turbulent velocity fluctuations over a geometric sequence of scales. Our goal is to study the convergence of solutions in the inviscid (more generally, vanishing regularization) limit and explain the universality of both the limiting solutions and the convergence process. We develop a renormalization group (RG) formalism representing this limit as dynamics in a space of flow maps. For the dyadic shell model, the RG dynamics has a fixed-point attractor, which determines universal limiting solutions. Deviations from the limiting solutions are also universal and given by a leading eigenmode (eigenvalue and eigenvector) of the linearized RG operator. Application to the Gledzer shell model reveals the RG attractor in the form of a closed invariant curve, while the Sabra shell model yields chaotic RG dynamics. An important consequence of the RG formalism is the understanding of the different roles of symmetry-preserving (canonical) and symmetry-breaking (e.g. viscous) regularizations.
@@ 1 Introduction
Our main result is the development of the renormalization group (RG) formalism. It describes the limit of vanishing regularization as the dynamics of an RG operator acting in the space of flow maps. A similar formalism was presented earlier for models on a discrete space-time lattice [38, 37, 36]. The extension to shell models in which time is continuous is carried out by introducing the concept of canonical (symmetry preserving) regularizations. For the dyadic model, the RG approach not only justifies the existence and universality of the limiting solutions as a fixed-point attractor of the RG dynamics, but also predicts new universal properties. This new universality is related to the leading eigenmode (eigenvalue and eigenvector) of the linearized RG operator, which determine deviations of regularized solutions from their limiting form. The usual viscous regularization is not canonical (it violates time scale invariance), but its RG analysis is mediated by auxiliary canonical regularizations. We also apply the RG approach to the Gledzer and Sabra shell models. In the former case, we show that the attractor of the RG dynamics is an invariant closed curve, which governes the vanishing regularization limit. For the Sabra model, the RG dynamics is chaotic with double exponential divergence of solutions.
In general, an RG formalism explores self-similarity of a system manifested at a large range of scales, but its precise form may vary a lot depending on the problem. In our case, this is the self-similarity of equations of motion corresponding to the ideal system. As we showed in [37], our RG approach has much in common with the Feigenbaum theory [17]; in particular, one can establish the analytical similarity for the model of digital turbulence. Opposite to the Kadanoff–Wilson RG approach [43] and its extension to fluid dynamics [19, 44, 4], we do not coarse-grain system properties but rather keep all details at small scales intact. One can also recognize in our formalism some features of the inverse RG [15, 21, 24], because our RG operator acts by adding an extra largest scale and reconstructing the ideal dynamics at that scale.
The paper is organized as follows. Section 2 introduces shell models. Section 3 defines the RG operator and relates it to a class of canonical regularizations. Section 4 studies the limit of vanishing regularization in terms of the fixed-point RG attractor for the dyadic model. Section 5 extends the results to the (non-canonical) viscous regularization. Section 6 studies the RG attractor in the Gledzer shell model. Section 7 describes the chaotic RG dynamics in the Sabra model. Section 8 summarizes the results and discusses further developments. Some technical derivations are gathered in the Appendix.
@@ 2 Ideal and viscous dyadic models
where the tildes denote a new solution for new initial and boundary conditions. It is useful to write Eq. (2.5) as
$\tilde{u}(t)=S^{+}u(t),\quad\tilde{a}=S^{+}a,\quad\tilde{b}(t)=\lambda u_{1}(t),$ (2.6)
where we define the scaling operators $S^{+}a=(\lambda a_{2},\lambda a_{3},\lambda a_{4},\ldots)$ and $S^{-}a=(0,a_{1}/\lambda,a_{2}/\lambda,\ldots)$. Similarly, the time scaling symmetry takes the form
$\tilde{u}(t)=\alpha u(\alpha t),\quad\tilde{a}=\alpha a,\quad\tilde{b}(t)=\alpha b(\alpha t),$ (2.7)
@@ 3 Canonical regularizations and RG operator
[…]
$\Phi_{t+s}(a,b)=\Phi_{s}(a^{\prime},b^{\prime}),\quad t,s\geq 0,$ (3.2)
where $a^{\prime}=\Phi_{t}(a,b)$ with the the time-shifted boundary condition $b^{\prime}(t^{\prime})=b(t+t^{\prime})$. This time shift takes into account non-autonomous dynamics due to time-dependent boundary condition.
@@
A flow map $\Phi$ is called renormalizable if, for any $a$ and $b$, there exists a unique solution $u_{1}(t)$ of the initial value problem
$\frac{du_{1}}{dt}=k_{1}f(b,u_{1},u_{2}),\quad u_{1}(0)=a_{1},$ (3.3)
where $u_{2}(t)=\tilde{u}_{1}(t)/\lambda$ with
$\tilde{u}(t)=\Phi_{t}(\tilde{a},\tilde{b}),\quad\tilde{a}=S^{+}a;\quad\tilde{b}(t)=\lambda u_{1}(t).$ (3.4)
Then, we introduce the RG operator $\Phi\mapsto\mathcal{R}[\Phi]$ acting on renormalizable flow maps as
$\mathcal{R}[\Phi]_{t}(a,b)=S^{-}\tilde{u}(t)+(u_{1}(t),0,0,\ldots)=(u_{1}(t),\tilde{u}_{1}(t)/\lambda,\tilde{u}_{2}(t)/\lambda,\ldots).$ (3.5)
@@
[…]
@@
From the definition of the scaling operators it is clear that $S^{-}S^{+}a=(0,a_{2},a_{3},\ldots)$, which differs from the identity in the first component. Hence, one can see that the relations (3.4) and (3.5) coincide with the scaling relation (2.6) for the components $n\geq 2$ of the solution $u(t)=\mathcal{R}[\Phi]_{t}(a,b)$. Since $\tilde{u}(t)$ is generated by the $N$-level regularization $\Phi$, it satisfies the ideal Eq. (2.1) for the shells $n=1,\ldots,N$. Then one can check that $u_{n}(t)$ satisfy the ideal Eq. (2.1) for the shells $n=2,\ldots,N+1$. The first relation in Eq. (3.3) provides the remaining ideal equation for $n=1$. ∎
For now, Definition 1 has an important flaw. Namely, combining the two symmetries (2.6) and (2.7), one can replace $\tilde{u}(t)$ in Eq. (3.4) by
$\tilde{u}(t)=\frac{1}{\alpha}\Phi_{t/\alpha}(\tilde{a},\tilde{b}),\quad\tilde{a}=\alpha S^{+}a;\quad\tilde{b}(t)=\alpha\lambda u_{1}(\alpha t).$ (3.6)
It is straightforward to check that Proposition 1 remains valid for this new formulation. This means that our definition of the RG operator is ambiguous. This ambiguity is eliminated by considering a specific subclass of canonical regularizations, as we show next.
@@
$\Phi_{t}(a,b)=\frac{1}{\alpha}\Phi_{t/\alpha}(\tilde{a},\tilde{b}),\quad\tilde{a}_{n}=\alpha a_{n},\quad\tilde{b}(t)=\alpha b(\alpha t),$ (3.7)
for any $a$, $b$ and $\alpha>0$. Clearly, this condition guarantees that $\tilde{u}(t)$ in Eq. (3.6) does not depend on $\alpha$.
@@
The important property of the RG operator $\mathcal{R}$ is that it is uniquely defined by the ideal model. Indeed, the ideal coupling function $f$ is the only model-dependent element in Definition 1. The space of canonical regularizations (the domain of $\mathcal{R}$) also depends on properties (symmetries) of the ideal model only. On the contrary, all the information on how solutions are regularized is contained in the flow map $\Phi$ itself. These features highlight the distinct roles of the ideal system and regularization in the RG dynamics: the first determines the RG operator $\mathcal{R}$, and the second the initial flow map $\Phi$.
@@
Finally, we observe that the viscous regularization (2.8) is not canonical, because the dissipative term is linear (not quadratic). This leads to important consequences that we investigate later in Section 5. The cutoff of models (3.8) mimics Large Eddy Simulation (LES) closures in fluid dynamics [40]: the dissipative term can be written as $k_{n}|u_{n}|u_{n}=\nu_{n}(u)k_{n}^{2}u_{n}$ with the effective eddy-viscosity $\nu_{n}(u)=|u_{n}|/k_{n}$.
@@ 4 Fixed-point RG attractor
An immediate consequence of the limit (4.2) is that the limiting flow map provides solutions $u^{\infty}(t)=\Phi^{\infty}_{t}(a,b)$ of the ideal IBVP. Moreover, these limiting solutions are universal: they do not depend on the choice of regularization, as long as the initial flow map $\Phi^{(0)}$ belongs to the basin of the attraction. Taking the limit $N\to\infty$ in both sides of Eq. (4.1), one can see that $\Phi^{\infty}$ satisfies the fixed-point condition $\Phi^{\infty}=\mathcal{R}[\Phi^{\infty}]$. This fixed-point condition is understood in the sense that Eqs. (3.3)–(3.5) of Definition 1 are satisfied for $\Phi=\Phi^{\infty}$ and $\mathcal{R}[\Phi]=\Phi^{\infty}$.
@@
[…]
@@
Numerical simulations of the regularized systems (3.8) are performed using the ode45 and ode15s solvers in Matlab with very high accuracy. Note that we do not need to solve the RG Eqs. (3.3)–(3.5) numerically. In fact, we already proved that they are satisfied by solutions of regularized systems (3.8). Therefore, the RG formalism is only used for interpreting the observed behavior in the inviscid limit.
@@ 5 RG approach to the viscous model
[…]
@@ 5.1 Auxiliary regularized model
Hence, the first-shell Eq. (5.3) takes the ideal form asymptotically for large $N$, i.e., it yields Eq. (3.3) of Definition 1. The rest of the argument is the same as for the regularized system (3.8) (see Section 3 and Appendix 9.2) and leads to the relation
@@ 5.2 Viscous model and inviscid limit
$u^{(\nu)}(t+dt)=\Phi^{(\nu)}_{dt}\left(u^{(\nu)}(t),b^{\prime}\right)=\Phi^{(N,\beta)}_{dt}\left(u^{(\nu)}(t),b^{\prime}\right),\quad\beta=\frac{\nu k_{N}}{|u_{N}^{(\nu)}(t)|},$ (5.9)
where $b^{\prime}(t^{\prime})=b(t+t^{\prime})$ is the time-shifted boundary condition. Let us select $N=N_{t}^{(\nu)}(a,b)$ as the largest shell number providing the parameter $\beta=\beta_{t}^{(\nu)}(a,b)=\nu k_{N}/|u_{N}^{(\nu)}(t)|\leq 1$. The time dependence of such $N$ and $\beta$ for different (decreasing) viscosities $\nu$ and specific initial and boundary conditions are shown Fig. 2(c,d). One can see that the inviscid limit $\nu\to 0$ implies $N\to\infty$. We argued in Section 5.1 that $\Phi^{(N,\beta)}\to\Phi^{\infty}$ as $N\to\infty$. Assuming that this convergence is uniform with respect to $\beta\leq 1$, the identity (5.9) implies that the evolution of the viscous solution can be approximated by the limiting flow map $\Phi^{\infty}$ at all times in the inviscid limit. This explains why the viscous regularization yields the same inviscid limit $\Phi^{(\nu)}\to\Phi^{\infty}$. This convergence is verified numerically in Fig. 2(a) by comparing solutions for $\nu=10^{-8}$ and $(N,\beta)=(20,1)$.
@@ 5.3 Asymptotic form of deviations
$u^{(\nu_{N})}(t+dt)\approx\Phi^{\infty}_{dt}\big{(}u^{(\nu_{N})}(t),b^{\prime}\big{)}+c_{\beta}\rho^{N}\Omega_{dt}\big{(}u^{(\nu_{N})}(t),b^{\prime}\big{)}.$ (5.11)
@@
$\Phi^{\infty}_{dt}\big{(}u^{(\nu_{N})}(t),b^{\prime}\big{)}\approx\Phi^{\infty}_{dt}\big{(}u^{\infty}(t),b^{\prime}\big{)}+\rho^{N}\delta\Phi^{\infty}_{dt}\big{(}v(t);u^{\infty}(t),b^{\prime}\big{)},$ (5.13)
where $\delta\Phi^{\infty}_{dt}(\delta u;u^{\infty}(t),b^{\prime})$ is the variational derivative of $\Phi^{\infty}_{dt}(u^{\infty}(t)+\delta u,b^{\prime})$. Similarly, the last term in Eq. (5.11) to the first-order becomes
$c_{\beta}\rho^{N}\Omega_{dt}\big{(}u^{(\nu_{N})}(t),b^{\prime}\big{)}\approx c_{\beta}\rho^{N}\Omega_{dt}\big{(}u^{\infty}(t),b^{\prime}\big{)}.$ (5.14)
Finally, using Eqs. (5.12)–(5.14) in the relation (5.11), cancelling the zero-oder terms $u^{\infty}(t+dt)=\Phi^{\infty}_{dt}\big{(}u^{\infty}(t),b^{\prime}\big{)}$ and then the common factor $\rho^{N}$, we obtain
$v(t+dt)\approx\delta\Phi^{\infty}_{dt}\left(v(t);u^{\infty}(t),b^{\prime}\right)+c_{\beta}\Omega_{dt}\big{(}u^{\infty}(t),b^{\prime}\big{)}.$ (5.15)
@@
6 Attracting closed invariant curve of RG dynamics
In this section, we further explore the dynamical systems approach to the vanishing regularization limit. We present here a more complex example of an RG attractor represented by a closed invariant curve. In this case regularized solutions do not converge in the limit $N\to\infty$, but one can describe them universally in terms of a one-parameter family of flow maps.